OpenStudy (anonymous):

2 . -23 = -2a - 10 + a ( 1point ) A. a = 31 b. a = 13 c. a = -11 d. a = 5

OpenStudy (gorv):

add 10 on both side

OpenStudy (gorv):

-23+10=-2a now divide both side by -2

OpenStudy (gorv):

-23+10=??

OpenStudy (anonymous):

13

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@Here_to_Help15

Nnesha (nnesha):

 combine like terms which is -2a + a = ???